# Streaming Operator Inference for Model Reduction of

### v1

@@ 3.3.1 Errors from Streaming SVD
$\|\mathbf{E}_{D}\|_{2}\leq\frac{\beta_{1}}{\sqrt{\min\{n,K\}}}\tau_{v},\qquad\|\mathbf{E}_{R}\|_{F}\leq\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v}.$ (18)
For the iSVD-LS/RLS paradigms (reformulation (15)),
$\|\mathbf{E}_{D}\|_{2}\leq\frac{\sigma_{1}\beta_{2}}{\sqrt{r}(\sigma_{1}+\varepsilon)}\tau_{w}+\frac{\beta_{2}}{\sigma_{1}+\varepsilon}\varepsilon,\qquad\|\mathbf{E}_{R}\|_{F}\leq\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{w}+\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}\varepsilon.$ (19)
Constants $\beta_{1},\beta_{2}>0$ depend on $\sigma_{1}$, bound of $\mathbf{U}$, and dimension $K$ where $\sigma_{1}$ denotes the largest singular value of $\mathbf{X}$.
Combining these lemmas yields the main result, bounding the operator error for each paradigm.
Theorem 3.3 (Operator error bounds).
Suppose $\boldsymbol{\Gamma}=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{d})$ with $\gamma_{j}>0$ for all $j$, and assume that $\sqrt{\gamma_{\min}}=\min_{j}\sqrt{\gamma_{j}}\gg\sigma_{\min}(\mathbf{D}),\sigma_{\min}(\mathbf{D}+\mathbf{E}_{D})$. Under the assumptions of Lemma 3.2, the operator error in Lemma 3.1 satisfies the following bounds. For the iSVD-Project-LS/RLS paradigms,
@@
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}\leq\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{v}\Bigl(1+\frac{\alpha\beta_{1}}{\sqrt{\gamma_{\min}}}\Bigr).$ (20)
For the iSVD-LS/RLS paradigms,
@@
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}\leq\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{w}\Bigl(1+\frac{\alpha\beta_{2}}{\sqrt{\gamma_{\min}}}\Bigr)+\frac{(\alpha\beta_{2}+\sqrt{\gamma_{\min}})\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}}{\gamma_{\min}}\varepsilon.$ (21)
@@
These theoretical results yield two key observations. First, when $\varepsilon_{w,j}>\varepsilon_{v,j}$ for all $j$, as is the case for iSVD algorithms satisfying (10), the iSVD-Project-LS/RLS paradigms can yield lower operator error than iSVD-LS/RLS. This follows because the bound (21) depends on $\tau_{w}$ and includes additional terms in $\varepsilon$, whereas the bound (20) depends only on $\tau_{v}$. Thus, projection paradigms are preferable when operator accuracy is critical and the additional computational cost and data passes are acceptable.
@@
Second, when the spectral decay is slow (e.g., the turbulent channel flow example in Section 4.3), the errors $\varepsilon_{w,j}$, $\varepsilon_{v,j}$, and $\varepsilon$ increase due to an insufficient reduced rank $r$ to capture the dominant modes. In such cases, increasing the regularization parameter $\gamma_{\min}$ can mitigate operator error by reducing sensitivity to these perturbations. For proofs of Lemmas 3.1, 3.2 and 3.3, see Appendix C.
@@ 3.3.2 Errors from Recursive Least Squares
Overall, the total operator error in the iSVD-LS and iSVD-Project-LS frameworks is due solely to errors from streaming SVD defined in Theorem 3.3. In contrast, the total operator error in iSVD-RLS and iSVD-Project-RLS frameworks is a combination of errors from streaming SVD defined in Theorem 3.3 and errors from RLS updates (22), in which the RLS solution approaches $\widetilde{\mathbf{O}}$ solved using matrices from streaming SVD with batch LS. In Section 4, we assess the streaming operator error to evaluate the convergence behavior of the RLS algorithms in practice.
@@ 3.4 Practical Considerations
To improve conditioning, Streaming OpInf applies preprocessing to snapshot data: mean-centering ensures the POD basis captures dominant variation modes Taira et al. [2017], while min-max normalization (e.g., scaling to $[0,1]$ or [1,1][-1,1]) prevents bias toward larger-magnitude variables in multi-physics systems McQuarrie et al. [2021], Kramer et al. [2024]. This is crucial for turbulent channel flow (Section 4.3), where streamwise velocity dominates. When making predictions, these transformations are inverted to return solutions in physical coordinates. Computing mean and min-max values requires an additional data pass, efficiently handled via batch processing and parallelization.
@@ 4.2.2 Evaluation Metrics
For iSVD assessment, we measure subspace alignment using subspace angles and approximation quality using relative projection errors, as in Section 4.1.2. For streaming operator learning, we track only the MR-SOE at each stream with $M=1$. Unlike the Burgers’ equation case, we omit state error metrics because KSE’s chaotic dynamics cause exponentially diverging trajectories from small perturbations, making point-wise error metrics less meaningful. Hence, it is crucial that we capture the statistical properties and attractor geometry of the chaotic system using Quantities of Interest (QoIs) that show dynamical invariants, i.e., invariants to initial conditions and individual trajectories.
@@ 4.3.3 Results & Discussion
Figure 11 displays the friction velocity $u_{\tau}$ evolution over time. The top panel shows both the original and predicted friction velocities normalized by the target mean friction velocity $\overline{u_{\tau}}$ computed by averaging the calculated friction velocities over time. Streaming OpInf show similar temporal fluctuations compared to the original data for both training and testing phases, capturing the characteristic variations arising from the passage of turbulent structures near the wall. The bottom panel presents the relative errors on a logarithmic scale, revealing errors consistently below $10^{-2}$ for the majority of snapshots, with occasional spikes reaching $10^{-5}$ that correspond to particularly strong instantaneous overlaps between the predicted and true friction velocities.
Figure 12 presents the wall-normal profile of the time-averaged streamwise velocity component. The logarithmic scale on the wall-normal coordinate axis emphasizes the near-wall region where velocity gradients are steepest Bredberg [2000]. For both training and testing data, Streaming OpInf predictions closely match the original data across the entire wall-normal extent, accurately capturing both the viscous sublayer near the wall and the turbulent/logarithmic layer in the outer region while adhering to the log-law (31Pinier et al. [2019], Marusic et al. [2013]. The relative errors, shown in the bottom panels, remain below $10^{-3}$ for training and $10^{-2}$ for testing throughout most of the domain, with slightly elevated errors in the near-wall region where the flow physics are most complex. The consistency between training and testing errors confirms the model’s ability to preserve statistical properties of the turbulent flow.
@@ Discussion
The results show that Streaming OpInf successfully learns accurate reduced models for a highly challenging large-scale turbulent flow problem. Starting from over 9 million degrees of freedom, the method achieves a state dimension reduction exceeding 31,000x while maintaining high predictive accuracy and enabling significantly faster state predictions. The streaming approach processes this large-scale dataset sequentially without requiring simultaneous access to all 8,000 snapshots, directly addressing the storage bottleneck in traditional batch methods.
@@ 5 Conclusion
Through comprehensive numerical experiments on benchmark problems such as the one-dimensional viscous Burgers’ equation, the Kuramoto-Sivashinsky equation, and a large-scale turbulent channel flow simulation, we have demonstrated accuracy on par with traditional batch OpInf methods while achieving over 99% reductions in memory usage. The experiments also highlighted the importance of practical considerations such as data standardization and regularization in enhancing model performance and stability.
@@
Tomoki Koike: Writing - review & editing, Writing - original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization; Prakash Mohan: Writing – review & editing, Supervision, Resources, Data curation; Marc T. Henry de Frahan: Writing - review & editing, Supervision, Resources, Data curation; Julie Bessac: Supervision, Funding Acquisition, Project administration; Elizabeth Qian: Conceptualization, Writing - review & editing, Supervision, Funding Acquisition, Project administration.
@@
Data and experiment code are available at https://github.com/smallpondtom/StreamingOpInf for reproducibility. The main implementations for Streaming OpInf are available at https://github.com/smallpondtom/LiftAndLearn.jl.
@@ Acknowledgement
The authors were supported by the Department of Energy Office of Science Advanced Scientific Computing Research, DOE Award DE-SC0024721. This work was authored in part by NLR for the U.S. Department of Energy (DOE), operated under Contract No. DE-AC36-08GO28308. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes. A portion of the research was performed using computational resources sponsored by the Department of Energy and located at the National Laboratory of the Rockies.
@@ Appendix B Recursive Least-Squares Algorithms
Appendix C Proof of Theorem 3.3
This section provides the proof of Theorem 3.3, which establishes error bounds for operators learned via streaming SVD within the proposed Streaming OpInf paradigms. The theorem quantifies how errors in the streaming SVD, relative to the batch SVD, propagate through the data matrices $\mathbf{D}$ and $\mathbf{R}$ into the learned operators. We first prove the supporting Lemmas 3.1 and 3.2, then combine these results to establish the theorem.
C.1 Proof of Lemma 3.1
The proof relies on the following classical perturbation result for pseudoinverses.
@@
Let the data matrices $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ be defined as in (7), and define the perturbed data matrices as $\widetilde{\mathbf{D}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}$ and $\widetilde{\mathbf{R}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}$. The corresponding LS solutions are $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ and $\widetilde{\mathbf{O}}=\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}$, respectively. Adding and subtracting $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\widetilde{\mathbf{R}}$ yields
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}=\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}\|_{F}=\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\widetilde{\mathbf{R}}+\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\widetilde{\mathbf{R}}-\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}\|_{F}\leq\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{R}})\|_{F}+\|(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}-\widetilde{\mathbf{D}}^{\dagger})\widetilde{\mathbf{R}}\|_{F},$
where the inequality follows from the triangle inequality. Applying submultiplicativity of the mixed matrix norms Fang et al. [1994] to each term gives
@@
Applying Lemma C.1 to the second term and substituting $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{R}}\|_{F}=\|\mathbf{E}_{R}\|_{F}$ and $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{D}}\|_{2}=\|\mathbf{E}_{D}\|_{2}$ completes the proof. Note that $\|[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}\|_{F}=\|\mathbf{E}_{R}\|_{F}$ and $\|[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}\|_{2}=\|\mathbf{E}_{D}\|_{2}$. ∎
C.2 Proof of Lemma 3.2
Before proceeding, we introduce notation for the full-state data matrix
$\mathbb{D}=\begin{bmatrix}\mathbf{X}^{\top}&(\mathbf{X}\odot\mathbf{X})^{\top}&\mathbf{U}^{\top}&\mathbf{1}_{K}\end{bmatrix}\in\mathbb{R}^{K\times(n+n^{2}+m+1)}.$
For the projection-based paradigms, the data matrices $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\widetilde{\mathbf{D}}$ can be expressed as
$\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu=[(\mathbb{D}\mathbb{V}_{r})^{\top\kern-0.5pt},\,\boldsymbol{\Gamma}^{1/2}]^{\top\kern-0.5pt}\quad\text{and}\quad\widetilde{\mathbf{D}}=[(\mathbb{D}\widetilde{\mathbb{V}}_{r})^{\top\kern-0.5pt},\,\boldsymbol{\Gamma}^{1/2}]^{\top\kern-0.5pt},$
where
$\mathbb{V}_{r}=\begin{bmatrix}\mathbf{V}_{r}&0&0&0\\ 0&\mathbf{V}_{r}\otimes\mathbf{V}_{r}&0&0\\ 0&0&\mathbf{I}_{m}&0\\ 0&0&0&1\end{bmatrix},\quad\widetilde{\mathbb{V}}_{r}=\begin{bmatrix}\widetilde{\mathbf{V}}_{r}&0&0&0\\ 0&\widetilde{\mathbf{V}}_{r}\otimes\widetilde{\mathbf{V}}_{r}&0&0\\ 0&0&\mathbf{I}_{m}&0\\ 0&0&0&1\end{bmatrix}$.
The matrices $\mathbb{V}_{r},\widetilde{\mathbb{V}}_{r}\in\mathbb{R}^{(n+n^{2}+m+1)\times(r+r^{2}+m+1)}$ have orthonormal columns.
Proof of Lemma 3.2.
The bound $|\sigma_{j}-\widetilde{\sigma}_{j}|\leq\varepsilon$ for all $j=1,\ldots,r$ follows directly from Weyl’s inequality [Golub and Van Loan, 2013, Cor. 8.6.2] under the assumption $\|\mathbf{X}-\widetilde{\mathbf{X}}\|_{2}\leq\varepsilon$.
Projection paradigms (iSVD-Project-LS/RLS)
We first bound the perturbation in the data matrix $\mathbf{D}$. By submultiplicativity,

### v2

@@ 3.3.1 Errors from Streaming SVD
$\|\mathbf{E}_{D}\|_{2}\leq\frac{\beta_{1}}{\sqrt{\min\{n,K\}}}\tau_{v},\qquad\|\mathbf{E}_{R}\|_{F}\leq\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{v},$ (17)
and the operator error between the batch solution $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ and the streaming solution $\widetilde{\mathbf{O}}=\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}$ satisfies
@@
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}\leq\frac{\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}}{\sqrt{\gamma_{\min}}}\tau_{v}\Bigl(1+\frac{\alpha\beta_{1}}{\sqrt{\gamma_{\min}}}\Bigr),$ (18)
where $\beta_{1}=2(\sigma_{1}^{4}+\sigma_{1}^{2}+\eta^{2}+K)^{1/2}\sqrt{\min\{n,K\}}$ with $\eta=\|\mathbf{U}\|_{2}$ and $\sigma_{1}$ is the largest singular value of $\mathbf{X}$.
Theorem 3.3 (Operator error bounds: Reformulation paradigms).
Under the setup of Theorem 3.2, define the right singular vector error $\tau_{w}=\|\mathbf{W}_{r}-\widetilde{\mathbf{W}}_{r}\|_{F}$. Then for the iSVD-LS/RLS paradigms, the data matrix perturbations satisfy
@@
$\|\mathbf{E}_{D}\|_{2}\leq\frac{\sigma_{1}\beta_{2}}{\sqrt{r}(\sigma_{1}+\varepsilon)}\tau_{w}+\frac{\beta_{2}}{\sigma_{1}+\varepsilon}\varepsilon,\qquad\|\mathbf{E}_{R}\|_{F}\leq\sigma_{1}\|\boldsymbol{\Delta}_{K}\|_{2}\tau_{w}+\sqrt{r}\|\boldsymbol{\Delta}_{K}\|_{2}\varepsilon,$ (19)
and the operator error satisfies
where $\beta_{2}=\sqrt{r}(\sigma_{1}+\varepsilon)\sqrt{1+(2\sigma_{1}+\varepsilon)^{2}}$ and $\sigma_{1}$ is the largest singular value of $\mathbf{X}$.
@@
These theoretical results yield two key observations. First, when $\tau_{w}>\tau_{v}$, as is the case for iSVD algorithms satisfying (10), the iSVD-Project-LS/RLS paradigms can yield lower operator error than iSVD-LS/RLS. This follows because the bound (20) depends on $\tau_{w}$ and includes additional terms in $\varepsilon$, whereas the bound (18) depends only on $\tau_{v}$. Thus, projection paradigms are preferable when operator accuracy is critical and the additional computational cost and data passes are acceptable.
@@
Second, when the spectral decay is slow (e.g., the turbulent channel flow example in Section 4.3), the errors $\tau_{w}$, $\tau_{v}$, and $\varepsilon$ increase due to an insufficient reduced rank $r$ to capture the dominant modes. In such cases, increasing the regularization parameter $\gamma_{\min}$ can mitigate operator error by reducing sensitivity to these perturbations.
@@ 3.3.2 Errors from Recursive Least Squares
Overall, the total operator error in the iSVD-LS and iSVD-Project-LS frameworks is due solely to errors from streaming SVD defined in Theorems 3.2 and 3.3. In contrast, the total operator error in iSVD-RLS and iSVD-Project-RLS frameworks is a combination of operator errors in Theorems 3.2 and 3.3 and errors from RLS updates (21), in which the RLS solution approaches $\widetilde{\mathbf{O}}$ solved using matrices from streaming SVD with batch LS. In Section 4, we assess the streaming operator error to evaluate the convergence behavior of the RLS algorithms in practice.
@@ 3.4 Practical Considerations
To improve conditioning, we recommend preprocessing of snapshot data: mean-centering ensures the POD basis captures dominant variation modes Taira et al. [2017], while min-max normalization (e.g., scaling to $[0,1]$ or [1,1][-1,1]) prevents bias toward larger-magnitude variables in multi-physics systems McQuarrie et al. [2021], Kramer et al. [2024]. This is crucial for turbulent channel flow (Section 4.3), where streamwise velocity dominates. When making predictions, these transformations are inverted to return solutions in physical coordinates. Computing mean and min-max values requires an additional data pass, efficiently handled via batch processing and parallelization.
@@ 4.2.2 Evaluation Metrics
For iSVD assessment, we measure subspace alignment using subspace angles and approximation quality using relative projection errors, as in Section 4.1.2. For streaming operator learning, we track only the MR-SOE at each stream with $M=1$. Unlike the Burgers’ equation case, we omit state error metrics because KSE’s chaotic dynamics cause exponentially diverging trajectories from small perturbations, making point-wise error metrics less meaningful. Hence, it is crucial that we capture the statistical properties and attractor geometry of the chaotic system using Quantities of Interest (QoIs) that show dynamical invariants, i.e., invariants to initial conditions and individual trajectories. This is also true for the large-scale turbulent flow example considered in Section 4.3.
@@ 4.3.3 Results & Discussion
Figure 11 displays the friction velocity $u_{\tau}$ evolution over time. The top panel shows both the original and predicted friction velocities normalized by the target mean friction velocity $\overline{u_{\tau}}$ computed by averaging the calculated friction velocities over time. Streaming OpInf show similar temporal fluctuations compared to the original data for both training and testing phases, capturing the characteristic variations arising from the passage of turbulent structures near the wall. The bottom panel presents the relative errors on a logarithmic scale, revealing errors consistently below $10^{-2}$ for the majority of snapshots, with occasional downward spikes decreasing to $10^{-5}$ that correspond to instantaneous overlaps between the predicted and true friction velocities.
Figure 12 presents the wall-normal profile of the time-averaged streamwise velocity component. The logarithmic scale on the wall-normal coordinate axis emphasizes the near-wall region where velocity gradients are steepest Bredberg [2000]. For both training and testing data, Streaming OpInf predictions closely match the original data across the entire wall-normal extent, accurately capturing both the viscous sublayer near the wall and the turbulent/logarithmic layer in the outer region while adhering to the log-law (30Pinier et al. [2019], Marusic et al. [2013]. The relative errors, shown in the bottom panels, remain below $10^{-3}$ for training and $10^{-2}$ for testing throughout most of the domain, with slightly elevated errors in the near-wall region where the flow physics are most complex. The consistency between training and testing errors confirms the model’s ability to preserve statistical properties of the turbulent flow.
@@ Discussion
The results demonstrate that Streaming OpInf enables accurate reduced model learning for a large-scale turbulent flow problem where batch OpInf is infeasible. With over 9 million degrees of freedom and 8,000 snapshots, the dataset exceeds practical memory limits for simultaneous storage and batch processing. By processing snapshots sequentially, Streaming OpInf makes this type of model reduction tractable, achieving a state dimension reduction exceeding 31,000x while maintaining high predictive accuracy and enabling significantly faster state predictions. This fundamentally expands the scope of OpInf to large-scale problems previously inaccessible to batch methods.
@@ 5 Conclusion
Through comprehensive numerical experiments on benchmark problems such as the one-dimensional viscous Burgers’ equation, the Kuramoto-Sivashinsky equation, and a large-scale turbulent channel flow simulation, we have achieved over 99% memory savings compared to batch OpInf—expanding the scope of OpInf to previously infeasible large-scale applications. The experiments also highlighted the importance of practical considerations such as data standardization and regularization in enhancing model performance and stability.
@@
Tomoki Koike: Writing - review & editing, Writing - original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization; Prakash Mohan: Writing – original draft, review & editing, Supervision, Resources, Data curation; Marc T. Henry de Frahan: Writing - review & editing, Supervision, Resources, Data curation; Julie Bessac: Supervision, Funding Acquisition, Project administration; Elizabeth Qian: Conceptualization, Writing - review & editing, Supervision, Funding Acquisition, Project administration.
@@
Data and experiment code are available at https://github.com/smallpondtom/StreamingOpInf for reproducibility.
@@ Acknowledgement
The authors were supported by the Department of Energy Office of Science Advanced Scientific Computing Research, DOE Award DE-SC0024721. This work was authored in part by NLR for the U.S. Department of Energy (DOE), operated under Contract No. DE-AC36-08GO28308. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes. A portion of the research was performed using computational resources sponsored by the Department of Energy’s Office of Critical Minerals and Energy Innovation and located at the National Laboratory of the Rockies.
@@ Appendix B Recursive Least-Squares Algorithms
Appendix C Proof of Operator Error Bounds
This section establishes the error bounds in Theorems 3.2 and 3.3, which quantify how errors from streaming SVD propagate through the data matrices into the learned operators. We begin by proving the general operator perturbation bound in Lemma 3.1, then derive the specific bounds for each streaming paradigm.
C.1 General Operator Perturbation Bound
The proof of Lemma 3.1 relies on a classical perturbation result for pseudoinverses.
@@ C.1 General Operator Perturbation Bound
Let $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu$ and $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ be the data matrices defined in (7), and let $\widetilde{\mathbf{D}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}$ and $\widetilde{\mathbf{R}}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu+[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}$ denote the perturbed data matrices. The corresponding LS solutions are $\mathbf{O}=\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu$ and $\widetilde{\mathbf{O}}=\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}$. Adding and subtracting $\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\widetilde{\mathbf{R}}$ and applying the triangle inequality yields
$\|\mathbf{O}-\widetilde{\mathbf{O}}\|_{F}$ $=\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{D}}^{\dagger}\widetilde{\mathbf{R}}\|_{F}$
[…]
$\leq\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{R}})\|_{F}+\|(\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}-\widetilde{\mathbf{D}}^{\dagger})\widetilde{\mathbf{R}}\|_{F}.$
Applying the submultiplicativity property $\|\mathbf{A}\mathbf{B}\|_{F}\leq\|\mathbf{A}\|_{2}\|\mathbf{B}\|_{F}$ to each term gives
@@
Since $\|[\mathbf{E}_{R}^{\top},\mathbf{0}_{r\times d}]^{\top\kern-0.5pt}\|_{F}=\|\mathbf{E}_{R}\|_{F}$ and $\|[\mathbf{E}_{D}^{\top},\mathbf{0}_{d\times d}]^{\top\kern-0.5pt}\|_{2}=\|\mathbf{E}_{D}\|_{2}$, we have $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{R}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{R}}\|_{F}=\|\mathbf{E}_{R}\|_{F}$ and $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu-\widetilde{\mathbf{D}}\|_{2}=\|\mathbf{E}_{D}\|_{2}$. Applying Lemma C.1 to bound $\|\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu^{\dagger}-\widetilde{\mathbf{D}}^{\dagger}\|_{2}$ completes the proof. ∎
C.2 Operator Error Bounds for Each Paradigm
We now prove Theorems 3.2 and 3.3 by deriving explicit bounds on the data matrix perturbations $\mathbf{E}_{D}$ and $\mathbf{E}_{R}$ for each streaming paradigm, then applying Lemma 3.1. We establish several auxiliary results needed for both proofs.
Lemma C.2 (Norm of the pseudoinverse).
Proof.
This follows directly from the SVD $\mathbf{A}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ and the pseudoinverse definition $\mathbf{A}^{\dagger}=\mathbf{V}\boldsymbol{\Sigma}^{\dagger}\mathbf{U}^{\top}$. ∎
The following result from Weyl’s inequality [Golub and Van Loan, 2013, Cor. 8.6.2] bounds the singular value perturbations.
Lemma C.3 (Singular value perturbation).
If $\|\mathbf{X}-\widetilde{\mathbf{X}}\|_{2}\leq\varepsilon$, then $|\sigma_{j}-\widetilde{\sigma}_{j}|\leq\varepsilon$ for all $j=1,\ldots,r$.
C.2.1 Proof of Theorem 3.2
Proof.
We derive bounds on the data matrix perturbations for the projection paradigms, then combine them with Lemma 3.1 to obtain the operator error bound.
Bounding the left-hand side perturbation $\mathbf{E}_{D}$
To systematically bound $\mathbf{E}_{D}$ for the projection paradigms, we introduce the full-state data matrix
$\mathbb{D}=\begin{bmatrix}\mathbf{X}^{\top}&(\mathbf{X}\odot\mathbf{X})^{\top}&\mathbf{U}^{\top}&\mathbf{1}_{K}\end{bmatrix}\in\mathbb{R}^{K\times(n+n^{2}+m+1)},$
and the projection matrices
$\mathbb{V}_{r}=\begin{bmatrix}\mathbf{V}_{r}&0&0&0\\ 0&\mathbf{V}_{r}\otimes\mathbf{V}_{r}&0&0\\ 0&0&\mathbf{I}_{m}&0\\ 0&0&0&1\end{bmatrix},\quad\widetilde{\mathbb{V}}_{r}=\begin{bmatrix}\widetilde{\mathbf{V}}_{r}&0&0&0\\ 0&\widetilde{\mathbf{V}}_{r}\otimes\widetilde{\mathbf{V}}_{r}&0&0\\ 0&0&\mathbf{I}_{m}&0\\ 0&0&0&1\end{bmatrix}$,
where $\mathbb{V}_{r},\widetilde{\mathbb{V}}_{r}\in\mathbb{R}^{(n+n^{2}+m+1)\times(r+r^{2}+m+1)}$ have orthonormal columns. This notation allows us to express the data matrices compactly as
$\mskip 1.5mu\overline{\mskip-1.5mu{\mathbf{D}}\mskip-1.5mu}\mskip 1.5mu=[(\mathbb{D}\mathbb{V}_{r})^{\top\kern-0.5pt},\,\boldsymbol{\Gamma}^{1/2}]^{\top\kern-0.5pt}\quad\text{and}\quad\widetilde{\mathbf{D}}=[(\mathbb{D}\widetilde{\mathbb{V}}_{r})^{\top\kern-0.5pt},\,\boldsymbol{\Gamma}^{1/2}]^{\top\kern-0.5pt}.$
By submultiplicativity,